MINDTAP Q Search this Homework (Ch 03) 3. Galns from trade Consider two neighboring island countries called Euphoria and Bellissima. They each have 4 million labor hours available per week that they can use to produce jeans, corn, or a combination of both. The following table shows the amount of jeans or corn that can be produced using 1 hour of labor. Jeans (Pairs per hour of labor) Country Euphoria Bellissima Corn (Bushels per hour of labor) 20 16 Initially, suppose Bellissima uses 1 million hours of labor per week to produce jeans and 3 million hours per week to produce corn, while Euphoria uses 3 million hours of labor per week to produce jeans and 1 million hours per week to produce corn. Consequently, Euphoria produces 15 million pairs of jeans and 20 million bushels of corn, and Bellissima produces 8 million pairs of jeans and 48 million bushels of corn. Assume there are no other countries willing to trade goods, so, in the absence of trade between these two countries, each country consumes the amount of jeans and corn it produces Euphorias opportunity cost of producing 1 pair of jeans is 4 bushelsof corm, and Bellissimas opportunity cost of producing 1 pair of jeans is 2 bushels of corn. Therefore, Bellissima has a comparative advantage in the production of jeans, and Euphoria has a comparative advantage in the production of corn.
Answer #1

Your all answers filled are correct.

Euphoria opportunity cost of 1 pair of jeans = 20/5 = 4 corn

Bellissima opportunity cost of 1 pair of jeans = 16/8 = 2 corn

Bellissima has a comparative advantage in jeans and euphoria has a comparative advantage in corn